OpenStudy (anonymous):

how do i solve this: ( 3/x-1 ) + 2/x =4

OpenStudy (matheducatormcg):

do you mean \[\frac{3}{x-1}+\frac{2}{x}=4\]

OpenStudy (anonymous):

yes

OpenStudy (matheducatormcg):

these types of equations can be solved by finding the LCD of all the fractions and multiplying each fraction by that LCD. The LCD will divide out the denominator leaving you a much easier equation. the LCD for these fractions is \[x(x-1)\]

OpenStudy (matheducatormcg):

so, \[x(x-1)\left( \frac{3}{x-1} \right)+x(x-1)\left( \frac{2}{x} \right)=4x(x-1)\]

OpenStudy (matheducatormcg):

you need to distribute and solve for x now. let me know if you need me to go further.

OpenStudy (anonymous):

so my new equation would be 2/x=4x-4 right?

OpenStudy (matheducatormcg):

not quite. the first fraction has the x-1 divide out resulting in 3x the next fraction has the x divide out resulting in 2(x-2), which is 2x-4 the other side of the equation becomes 4x^2-4x by the distributive property.

OpenStudy (matheducatormcg):

oops... second fraction divided out to 2(x-1), which is 2x-2. sorry about that

OpenStudy (matheducatormcg):

your new equation should look like \[3x+2x-2=4x^2-4x\]

OpenStudy (matheducatormcg):

solve by factoring or quadratic formula now.

OpenStudy (anonymous):

okay